### 08 Dec 2023 v1

# All-Sky Search for Transient Astrophysical Neutrino Emission with 10 Years of IceCube Cascade Events

Abstract
We present the results of a time-dependent search for neutrino flares in data collected by IceCube between May 2011 and 2021.
This data set contains cascade-like events originating from charged-current electron neutrino and tau neutrino interactions and all-flavor neutral-current interactions.
IceCube’s previous all-sky searches for neutrino flares used data sets consisting of track-like events originating from charged-current muon neutrino interactions.
The cascade data sets are statistically independent of the track data sets and provide a new opportunity to observe the transient all-sky landscape.
This search uses the spatial, temporal, and energy information of the cascade-like events to conduct searches for the most statistically significant neutrino flares in the northern and southern skies. No statistically significant time-dependent neutrino emission was observed. For the most statistically significant location in the northern sky, 0.71, and in the southern sky, 0.51. These results are compatible with the background hypothesis. Assuming an E spectrum from the diffuse astrophysical neutrino flux as measured with cascades, these results are used to calculate upper limits at the 90% confidence level on neutrino flares of varying duration and constrain the contribution of these flares to the diffuse astrophysical neutrino flux. These constraints are independent of a specified class of astrophysical objects and show that multiple unresolved transient sources may contribute to the diffuse astrophysical neutrino flux.
1 Introduction
The astrophysical processes responsible for producing, accelerating, and propagating high-energy cosmic rays have not been resolved. High-energy neutrinos can provide insights into the origins of cosmic rays. In 2013, IceCube reported observations of the diffuse astrophysical neutrino flux (Aartsen et al., 2013a, 2014a, 2015a). However, the majority of the diffuse astrophysical neutrino flux still has unresolved origins. Since then, IceCube has observed evidence of time-dependent and steady-state neutrino emission from astrophysical objects. In 2017, IceCube detected a 290 TeV neutrino (IceCube-170922A) in spatial coincidence with the blazar TXS 0506+056; IceCube-170922A was also temporally coincident with enhanced multi-wavelength activity in the blazar (Aartsen et al., 2018a). The statistical significance of this spatial and temporal coincidence was reported at 3. An archival search of 9.5 years of IceCube data found further evidence at a statistical significance of 3.5 for a neutrino flare that occurred between September 2014 and March 2015 that was not coincident with gamma-ray emission (Aartsen et al., 2018a, b). IceCube observed evidence for continuous neutrino emission from the Seyfert galaxy NGC 1068 with a significance of 4.2 (Abbasi et al., 2022). Recently, IceCube has also observed steady diffuse neutrino emission from the Milky Way at a significance of 4.5 (Abbasi et al., 2023). In the past decade, time-domain multi-messenger astronomy has observed multiple breakthroughs that serve potential insight into which sources are capable of contributing to the diffuse astrophysical neutrino flux. Not all astrophysical transient events are expected to produce a neutrino flux; however, hadronic astrophysical transient events are primary candidates for contributing to the diffuse neutrino flux (Murase & Bartos, 2019). In order to remain model-independent, this analysis excludes multi-messenger information and solely uses 10 years of IceCube’s cascade data to perform a time-dependent search for neutrino flares across the entire sky.
The IceCube Neutrino Observatory is located at the geographic South Pole and occupies a cubic kilometer of ice instrumented with 5,160 digital optical modules (DOMs). The DOMs are situated on 86 readout and support cables that have been frozen into the Antarctic glacier at depths between 1450 m – 2450 m (Abbasi et al., 2009; Aartsen et al., 2017). The primary in-ice DOMs array is composed of 78 readout and support cables; the DOMs have a vertical spacing of 17 m and a horizontal spacing of 125 m. This design allows IceCube to detect the Cherenkov radiation from charged particles that are created by neutrinos with energies from 100 GeV – 10 PeV interacting with the Antarctic ice. The remaining 8 cables were used to create a higher density sub-volume of IceCube in order to detect neutrinos down to 10 GeV (Abbasi et al., 2012). Since 2011, IceCube has been fully operational and taking data with 99% uptime (Aartsen et al., 2017).
2 Search for Neutrino Flares
IceCube detects electron neutrinos (), muon neutrinos (), and tau neutrinos (), as well as their antiparticles. At Earth, we expect a flavor ratio ( : : ) of approximately 1:1:1 due to neutrino oscillations occurring over astronomical distances (Learned & Pakvasa, 1995). Depending on the interaction type, a neutrino will primarily produce two event topologies: cascades and tracks (Aartsen et al., 2014b). Cascade events are produced by all-flavor neutral-current interactions, as well as charged-current and interactions, which can produce hadronic and electromagnetic showers that typically have radial extents less than 20 m (Aartsen et al., 2014b). Due to the spatial extent of the showers and the scattering length of light within the Antarctic ice, which is small compared to the DOM spacing, cascade events appear nearly symmetric and have reconstructed angular resolutions of 5 – 15 depending on the energy of the neutrino (Aartsen et al., 2013b; Abbasi et al., 2023). Conversely, tracks are produced when interact with the ice through the charged-current channel and produce a muon that can travel for several kilometers while emitting Cherenkov light. At high energies, tracks have reconstructed angular resolutions 1 (Abbasi et al., 2021b). However, tracks have higher background rates due to the atmospheric muon and background in IceCube, whereas cascades have lower background rates due their distinct event topology.
IceCube has conducted multiple “untriggered” all-sky transient analyses (Abbasi et al., 2011, 2021c, 2021a), meaning we do not use external multi-messenger information to search for time-dependent neutrino emission. Untriggered searches allow IceCube to be sensitive to astrophysical phenomena that solely produce neutrinos or produce neutrinos and other astrophysical messengers at offset times. When searching for neutrino flares that have durations of days or less, this method is more sensitive than IceCube’s time-integrated searches (Braun et al., 2010). IceCube’s previous untriggered analyses only used track events due to their high event rate and small angular uncertainties. These analyses did not observe statistically significant time-dependent neutrino emission. Here, we use 10 years of cascade events to search for the most statistically significant spatial and temporal clustering of events in IceCube’s northern ( -5) and southern ( -5) sky, which are defined to be consistent with IceCube’s previous analyses. Additionally, the points within 10 of the celestial poles are excluded due to their low statistics and limited background estimation. The analysis of cascade events provides a new opportunity to use a statistically independent data set to observe the transient sky.
2.1 Cascade Data set
This analysis uses IceCube’s most recent cascade data set and was motivated by IceCube’s previous 7-year cascade data set (Aartsen et al., 2019; Abbasi et al., 2023). This cascade data set extends the previous data set by using novel event reconstruction and event selection techniques with neural networks and boosted decision trees. The new reconstruction improves IceCube’s sensitivity to astrophysical sources that emit at hard and soft spectra while obtaining an event rate that is an order of magnitude higher than the previous cascade data set (Abbasi et al., 2021d). The event selection procedures are able to efficiently reject atmospheric muons, leading to a data set that is dominated by neutrinos (Abbasi et al., 2023). Approximately 87% of the cascade events are estimated to be atmospheric neutrinos, 7% are astrophysical neutrinos, and the remainder being atmospheric muons (Abbasi et al., 2023). The atmospheric neutrino spectrum is comparatively softer than that of astrophysical neutrinos (Aartsen et al., 2015b). Thus, at energies between 10 TeV – 100 TeV, atmospheric neutrinos become the subdominant component of the cascade data set (Abbasi et al., 2023).
This data set was initially designed to search for time-integrated all-sky and Galactic diffuse emission. The cascade data set contains 59,592 cascade-like events detected with IceCube’s complete 86-string configuration from May 13, 2011 to May 27, 2021 with energies between 500 GeV – 5.35 PeV. The angular resolution of the data set is energy-dependent. At 1 TeV, the median angular resolution is 15; at 100 TeV, the median angular resolution is 10 and improves to 5 at PeV energies.
2.2 Analysis Methods
The clustering search uses a maximum likelihood method to identify time-dependent neutrino emission from point sources, similar to IceCube’s previous analyses (Braun et al., 2010; Abbasi et al., 2021c, a). The likelihood,
is comprised of spatial, energy, and temporal probability density functions (PDFs) which identify how signal-like or background-like a given cascade event is. To search for neutrino emission, the sky is partitioned into 12,288 pixels of equal solid angle using HEALPix111https://healpix.jpl.nasa.gov/ (Gorski et al., 2005). The spatial PDF accounts for the angular distance between the center of each pixel and the reconstructed arrival direction of each cascade event. Due to the rotation of the Earth, the angular distribution of background events is assumed to be uniform in right ascension. We also assume a Gaussian PDF to test for temporal clustering between given cascade events and and treat IceCube’s background event rate as uniform in time while also accounting for detector livetime. IceCube’s previous searches have been performed with both a Gaussian and a “box-shaped” temporal PDF which follows a Heaviside step function. Here we use the Gaussian PDF since it is less computationally intensive. The energy PDFs account for the astrophysical spectra and the energy distribution of the cascade events as a function of declination. We assume that for the signal hypothesis, the neutrino flux follows an unbroken power-law, E, where is the spectral index.
In this analysis, we define the null hypothesis such that there is no spatial and temporal neutrino clustering for a given area of the sky. The alternative hypothesis fits for the following four parameters at every pixel: number of signal events (), signal spectral index (), mean time of the flare ( [Modified Julian Date (MJD)]), and half-width flare duration ( [days]) in order to search for neutrinos that are spatially and temporally clustered. The alternative hypothesis indicates an excess of signal-like events that surpass the expected background. To test the null and alternate hypothesis we use the likelihood ratio test, as described in Wilks (1938), to construct a test statistic (TS),
In Equation 2, and refer to the livetime bounds of the cascade data set. The TS includes a marginalization term for the flare duration in addition to the likelihood ratio; this prevents a bias that arises from over-fitting for with shorter durations. IceCube’s previous untriggered flare searches have searched for flares by imposing a signal-over-background () threshold to reduce the computational complexity of the analysis. The threshold is used to seed potential flares with events that have a high probability of being a signal, based off their reconstructed direction and energy. Thus, only accounts for the spatial and energy PDFs of a given event. Previous all-sky track analyses had data sets with high data rates; these analyses used thresholds of to efficiently search for flares. Since the cascade data set has an order of magnitude fewer events than IceCube’s track data set, we relax this threshold to .
To calculate the TS, the events’ arrival directions are uniformly randomized in right ascension; the time of the event is also randomized. A TS is then calculated for each pixel in the all-sky search. This process is repeated – times in order to obtain an ensemble of TS that follows the null hypothesis. Each TS in the ensemble represents a different realization of background data. To minimize the computational intensiveness of building an ensemble of TS for each pixel, and to account for the declination dependence of the data, we build an ensemble of TS in 81 declination bins and one right ascension bin. The declination bins range from [-80, 80] and are evenly spaced in 2 increments. The TS for the data, TS, is then calculated for every pixel and compared to the corresponding ensemble of TS in order to calculate a local p-value (), which is the probability of getting a TS TS from a random background realization of our data. We report two local p-values, , for the data that respectively correspond to the most statistically significant pixels in the northern and southern sky. Since we test each pixel on the sky, the two for the northern and southern sky are corrected for statistical trials to report global p-values (). Considering randomized sky scans that correspond to different realizations of background data, we take the most statistically significant in the northern and southern sky from each sky scan in order to obtain respective distributions of the most statistically significant . For this analysis, is then the probability of observing a that is at least as significant as .
3 Results & Implications
No significant spatial and temporal clustering of astrophysical neutrinos was observed in this analysis. As a result, we calculated the 90% upper limits (U.L.) on neutrino fluxes (Figure 2 – left) for the all-sky hottest spot and we discuss implications for the observed diffuse astrophysical flux.
3.1 Time-Dependent All-Sky Scan
Figure 1 shows the experimental results of the time-dependent all-sky scan; it is centered on the most statistically significant pixel in the sky. In the northern sky, the most significant point is located at with a corresponding and a significance of The most significant point in the southern sky is located at with a respective and a significance of The “hottest” spots are separated by in declination and are located in the same extended “warm” spot. This is due in part to the angular resolution of the cascade events, the sensitivity of this data set, and the analysis methods, which separate the northern and southern skies at = -5 (Abbasi et al., 2023). Since cascade events are correlated with different pixels, the statistical fluctuations near the hottest spot have similar . Figure 1 shows that the hottest spots have consistent temporal best-fit parameters. After correcting the most statistically significant for the scan over every pixel in the sky, 0.71 in the northern sky, and 0.51 in the southern sky. The results of this all-sky search for transient astrophysical neutrino emission are consistent with the background-only hypothesis.
3.2 Diffuse Flux Constraints
IceCube has measured the diffuse astrophysical neutrino spectrum with cascade event topologies (Aartsen et al., 2020). Assuming a power-law of E from the measured diffuse cascade flux, we calculate the time-integrated fluxes of the sensitivity of this analysis and the resulting U.L. from data, each at the 90% confidence level (C.L.), at the location of the most significant point. To do so, TS and TS are used as the benchmark TS values at which the injected time-integrated flux levels would exceed these benchmarks 90% of the time. This is repeated for each . Plotted in Figure 2 – left are the time-integrated fluxes converted to time-averaged fluxes by distributing them over the entire livetime of the data set in order to compare to the measured diffuse cascade flux. As the flare duration increases, the maximum possible contribution from a single neutrino flare to the diffuse astrophysical neutrino flux increases. However, at least 100 neutrino flares of = 100 days are needed to account for the diffuse astrophysical flux.
The sensitivity and discovery potential time-integrated fluxes are shown as a function of in Figure 2 – right assuming an E spectrum. The discovery potential is calculated as the time-integrated flux in which corresponding TS values will exceed the background TS ensemble’s 5 threshold 90% of the time. IceCube’s signal acceptance and background rates differ as a function of declination, resulting in different sensitivity and discovery potential fluxes at different declinations. In both Figure 2 – left and right, as increases the background rates also increase, thus the corresponding time-integrated flux needed increases too.
4 conclusions
In the all-sky search for transient neutrino emission, no significant spatial and temporal clustering was observed. We provide upper limits on the time-averaged neutrino flux at the 90% confidence level for various assumed flare durations, as shown in Figure 2 – left. These all-sky upper limits show that depending on the flare duration hundreds or thousands of individual neutrino flares are needed in order to comprise the diffuse astrophysical neutrino flux. These all-sky upper limits are not dependent on a specified class of astrophysical transient objects; thus, multiple unresolved transient objects may contribute to the diffuse neutrino flux.
Cascades provide upper limits that can help constrain future searches for transient neutrino emission. Cascade data sets are considered to be independent of IceCube’s track data sets, and future searches for transient neutrino emission could combine tracks and cascades. In addition, the next generation of IceCube, IceCube-Gen2, will have the capability to conduct a more sensitive search for transient neutrino emission across the entire sky (Aartsen et al., 2021). This will provide us with the opportunity to observe the origin of high-energy astrophysical neutrinos and further motivate time-domain multi-messenger astronomy.
Acknowledgements
The IceCube collaboration acknowledges the significant contributions to this manuscript from Michael Kovacevich. USA – U.S. National Science Foundation-Office of Polar Programs, U.S. National Science Foundation-Physics Division, U.S. National Science Foundation-EPSCoR, U.S. National Science Foundation-Office of Advanced Cyberinfrastructure, Wisconsin Alumni Research Foundation, Center for High Throughput Computing (CHTC) at the University of Wisconsin–Madison, Open Science Grid (OSG), Partnership to Advance Throughput Computing (PATh), Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS), Frontera computing project at the Texas Advanced Computing Center, U.S. Department of Energy-National Energy Research Scientific Computing Center, Particle astrophysics research computing center at the University of Maryland, Institute for Cyber-Enabled Research at Michigan State University, Astroparticle physics computational facility at Marquette University, NVIDIA Corporation, and Google Cloud Platform; Belgium – Funds for Scientific Research (FRS-FNRS and FWO), FWO Odysseus and Big Science programmes, and Belgian Federal Science Policy Office (Belspo); Germany – Bundesministerium für Bildung und Forschung (BMBF), Deutsche Forschungsgemeinschaft (DFG), Helmholtz Alliance for Astroparticle Physics (HAP), Initiative and Networking Fund of the Helmholtz Association, Deutsches Elektronen Synchrotron (DESY), and High Performance Computing cluster of the RWTH Aachen; Sweden – Swedish Research Council, Swedish Polar Research Secretariat, Swedish National Infrastructure for Computing (SNIC), and Knut and Alice Wallenberg Foundation; European Union – EGI Advanced Computing for research; Australia – Australian Research Council; Canada – Natural Sciences and Engineering Research Council of Canada, Calcul Québec, Compute Ontario, Canada Foundation for Innovation, WestGrid, and Digital Research Alliance of Canada; Denmark – Villum Fonden, Carlsberg Foundation, and European Commission; New Zealand – Marsden Fund; Japan – Japan Society for Promotion of Science (JSPS) and Institute for Global Prominent Research (IGPR) of Chiba University; Korea – National Research Foundation of Korea (NRF); Switzerland – Swiss National Science Foundation (SNSF).